# Critical behavior of a programmable time-crystal lattice gas

### v1

@@ III Monte Carlo simulations of the time crystal lattice gas
As discussed in the previous section, a good order parameter for the reported phase transition is the average magnitude of the packing order parameter, $\langle|z_{m}|\rangle$, which is a measure of the particles’ condensation around the $m$ emergent localization centers. Figs. 3.(a)-(c) show the measured $\langle|z_{m}|\rangle$ as a function of the packing field coupling $\lambda$ for $\rho_{0}=1/3$, $m=1,~{}2,~{}3$ and different system sizes $L$. As expected, the order parameter remains small for $\lambda<\lambda_{c}^{(m)}$, approaching zero in the thermodynamic $L\to\infty$ limit. However, upon crossing $\lambda_{c}^{(m)}$, the value of $\langle|z_{m}|\rangle$ increases sharply but in a continuous way, a behavior typical of second-order continuous phase transitions. Remarkably, the measured $\langle|z_{m}|\rangle(\lambda)$ agrees nicely with the hydrodynamic predictions for the order parameter $\forall m$ even for moderate system sizes, see red dashed line in Figs. 3(a)-(c). We also measured the susceptibility across the transition as captured by the order parameter scaled fluctuations, $\langle\Delta|z_{m}|^{2}\rangle\equiv L\left(\langle|z_{m}|^{2}\rangle-\langle|z_{m}|\rangle^{2}\right)$, see Figs. 3.(d)-(f). The susceptibility $\langle\Delta|z_{m}|^{2}\rangle$ exhibits a steep peak around $\lambda_{c}^{(m)}$, the sharper the larger the system size $L$, strongly suggesting a divergence of the scaled order parameter fluctuations at the critical point $\lambda_{c}^{(m)}$ in the thermodynamic limit, as expected for a continuous phase transition. Note that, for each $L$, the finite-size peak in $\langle\Delta|z_{m}|^{2}\rangle$ is always located slightly after the critical point, $\lambda^{*}\gtrsim\lambda_{c}^{(m)}$, an observation in contrast with the behavior reported in the Kuramoto model of synchronization Daido (1990); Hong et al. (2015) and most probably related with the particle exclusion interactions that characterize the TCLG model. In order to determine the critical packing field coupling $\lambda_{c}^{(m)}$ from simulations, we also measured the Binder cumulant $U_{4}=1-\langle|z_{m}|^{4}\rangle/(3\langle|z_{m}|^{2}\rangle^{2})$ as a function of $\lambda$ Binder and Heermann (2010), see Figs. 3(g)-(i). It is well known that, due to the finite-size scaling properties of this observable, the curves $U_{4}(\lambda)$ measured for different values of $L$ should cross at an $L$-independent value of $\lambda$ that identifies the infinite-size critical point $\lambda_{c}^{(m)}$. This is indeed clearly observed in Figs. 3.(g)-(i), where we can also verify that these measured critical packing field couplings agree remarkably well with the hydrodynamic predictions for $\lambda_{c}^{(m)}$ $\forall m$ (vertical dashed lines), obtained from a local stability analysis of the homogeneous density solution, see Eq. (7).
@@
Figure 6: Invariance of condensate profiles. Collapse of the measured condensate density profiles for different packing orders $m=1,~{}2,~{}3$, for $\lambda=1.03\lambda_{c}^{(m)}$ (top panel) and $\lambda=2\lambda_{c}^{(m)}$ (bottom panel). The different markers correspond to simulation results with $L=1920$ and various $m$, while the dashed line corresponds to the scaled hydrodynamic solution. The driving field for each packing order is $\epsilon=2.5m$, and the global density is $\rho_{0}=1/3$.
The question remains as to whether the shape of the emergent particle condensates in the time-crystal phase is also independent of the packing order $m$ after some renormalization of parameters, as it is indeed predicted by hydrodynamics, see Eqs. (4)-(7). In particular, the prediction is that the $m$-th-order traveling-wave solution of Eq. (4) once time-translation symmetry is spontaneously broken corresponds to $m$ copies of the $m=1$ solution after an appropriate renormalization of driving parameters. This means that $\rho_{m}(\omega_{m}t-2\pi mx)=\rho_{1}(m\omega_{1}t-2\pi mx)$, where $\rho_{1}(\omega_{1}t-2\pi x)$ is a traveling-wave solution of Eq. (4) of velocity $\omega_{1}$ for $m=1$ and generic driving parameters $\epsilon_{1}$ and $\lambda_{1}$, while $\rho_{m}(\omega_{m}t-2\pi mx)$ is the corresponding traveling-wave solution of Eq. (4) of velocity $\omega_{m}=m\omega_{1}$ for arbitrary $m>1$ and driving parameters $\epsilon_{m}=m\epsilon_{1}$ and $\lambda_{m}=m\lambda_{1}$. To address this question in simulations, we measured in detail the density profiles for different values of $\lambda>\lambda_{c}^{(m)}$ deep into the time-crystal phase for $m=1,~{}2,~{}3$ and parameters $\epsilon$ and $\lambda$ scaled as described above, see Fig. 6. In agreement with the hydrodynamic theory, the measured condensate profiles match and collapse neatly on top of each other, $\rho_{m}(x)=\rho_{1}(mx)$, once the spatial coordinate $x$ is stretched by the corresponding packing order $m$. Moreover, the collapsed shape agrees accurately with the hydrodynamic curve predicted by the traveling-wave solution, see dashed lines in Fig. 6.

### v2

@@ III Monte Carlo simulations of the time crystal lattice gas
As discussed in the previous section, a good order parameter for the reported phase transition is the average magnitude of the packing order parameter, $\langle|z_{m}|\rangle$, which is a measure of the particles’ condensation around the $m$ emergent localization centers. Figs. 3.(a)-(c) show the measured $\langle|z_{m}|\rangle$ as a function of the packing field coupling $\lambda$ for $\rho_{0}=1/3$, m=1,2,3m=1,~2,~3 and different system sizes $L$. As expected, the order parameter remains small for $\lambda<\lambda_{c}^{(m)}$, approaching zero in the thermodynamic $L\to\infty$ limit. However, upon crossing $\lambda_{c}^{(m)}$, the value of $\langle|z_{m}|\rangle$ increases sharply but in a continuous way, a behavior typical of second-order continuous phase transitions. Remarkably, the measured $\langle|z_{m}|\rangle(\lambda)$ agrees nicely with the hydrodynamic predictions for the order parameter $\forall m$ even for moderate system sizes, see red dashed line in Figs. 3(a)-(c). We also measured the susceptibility across the transition as captured by the order parameter scaled fluctuations, $\langle\Delta|z_{m}|^{2}\rangle\equiv L\left(\langle|z_{m}|^{2}\rangle-\langle|z_{m}|\rangle^{2}\right)$, see Figs. 3.(d)-(f). The susceptibility $\langle\Delta|z_{m}|^{2}\rangle$ exhibits a steep peak around $\lambda_{c}^{(m)}$, the sharper the larger the system size $L$, strongly suggesting a divergence of the scaled order parameter fluctuations at the critical point $\lambda_{c}^{(m)}$ in the thermodynamic limit, as expected for a continuous phase transition. Note that, for each $L$, the finite-size peak in $\langle\Delta|z_{m}|^{2}\rangle$ is always located slightly after the critical point, $\lambda^{*}\gtrsim\lambda_{c}^{(m)}$, an observation in contrast with the behavior reported in the Kuramoto model of synchronization Daido (1990); Hong et al. (2015) and most probably related with the particle exclusion interactions that characterize the TCLG model. In order to determine the critical packing field coupling $\lambda_{c}^{(m)}$ from simulations, we also measured the Binder cumulant $U_{4}=1-\langle|z_{m}|^{4}\rangle/(3\langle|z_{m}|^{2}\rangle^{2})$ as a function of $\lambda$ Binder and Heermann (2010), see Figs. 3(g)-(i). It is well known that, due to the finite-size scaling properties of this observable, the curves $U_{4}(\lambda)$ measured for different values of $L$ should cross at an $L$-independent value of $\lambda$ that identifies the infinite-size critical point $\lambda_{c}^{(m)}$. This is indeed clearly observed in Figs. 3.(g)-(i), where we can also verify that these measured critical packing field couplings agree remarkably well with the hydrodynamic predictions for $\lambda_{c}^{(m)}$ $\forall m$ (vertical dashed lines), obtained from a local stability analysis of the homogeneous density solution, see Eq. (6).
@@
Figure 6: Invariance of condensate profiles. Collapse of the measured condensate density profiles for different packing orders m=1,2,3m=1,~2,~3, for $\lambda=1.03\lambda_{c}^{(m)}$ (top panel) and $\lambda=2\lambda_{c}^{(m)}$ (bottom panel). The different markers correspond to simulation results with $L=1920$ and various $m$, while the dashed line corresponds to the scaled hydrodynamic solution. The driving field for each packing order is $\epsilon=2.5m$, and the global density is $\rho_{0}=1/3$.
The question remains as to whether the shape of the emergent particle condensates in the time-crystal phase is also independent of the packing order $m$ after some renormalization of parameters, as it is indeed predicted by hydrodynamics, see Eqs. (4)-(6). In particular, the prediction is that the $m$-th-order traveling-wave solution of Eq. (4) once time-translation symmetry is spontaneously broken corresponds to $m$ copies of the $m=1$ solution after an appropriate renormalization of driving parameters. This means that $\rho_{m}(\omega_{m}t-2\pi mx)=\rho_{1}(m\omega_{1}t-2\pi mx)$, where $\rho_{1}(\omega_{1}t-2\pi x)$ is a traveling-wave solution of Eq. (4) of velocity $\omega_{1}$ for $m=1$ and generic driving parameters $\epsilon_{1}$ and $\lambda_{1}$, while $\rho_{m}(\omega_{m}t-2\pi mx)$ is the corresponding traveling-wave solution of Eq. (4) of velocity $\omega_{m}=m\omega_{1}$ for arbitrary $m>1$ and driving parameters $\epsilon_{m}=m\epsilon_{1}$ and $\lambda_{m}=m\lambda_{1}$. To address this question in simulations, we measured in detail the density profiles for different values of $\lambda>\lambda_{c}^{(m)}$ deep into the time-crystal phase for m=1,2,3m=1,~2,~3 and parameters $\epsilon$ and $\lambda$ scaled as described above, see Fig. 6. In agreement with the hydrodynamic theory, the measured condensate profiles match and collapse neatly on top of each other, $\rho_{m}(x)=\rho_{1}(mx)$, once the spatial coordinate $x$ is stretched by the corresponding packing order $m$. Moreover, the collapsed shape agrees accurately with the hydrodynamic curve predicted by the traveling-wave solution, see dashed lines in Fig. 6.
@@ IV Discussion
Further, the packing-field mechanism explored here could be experimentally tested using colloidal fluids confined in quasi-$1d$ circular optical traps Cereceda-López et al. (2023); Cereceda-López (2023); Cereceda-López et al. (2024) or channels Lutz et al. (2004); Villada-Balbuena et al. (2021). In particular, optical tweezers Kumar and Bechhoefer (2018); Grier (1997); Ortiz-Ambriz et al. (2014); Martínez et al. (2015, 2017); Rodrigo et al. (2018) could be employed to implement the packing field by dynamically adjusting forces based on the system configuration, via feedback loop protocols. Such experimental setups, already used to study nonequilibrium phase behavior in colloidal systems, offer a promising platform to realize and probe the time-crystal phases discussed in this work.